### 30 Nov 2023 v1

# Large Language Models for Travel Behavior Prediction

Abstract
Travel behavior prediction is a fundamental task in transportation demand management. The conventional methods for travel behavior prediction rely on numerical data to construct mathematical models and calibrate model parameters to represent human preferences. Recent advancement in large language models (LLMs) has shown great reasoning abilities to solve complex problems. In this study, we propose to use LLMs to predict travel behavior with prompt engineering without data-based parameter learning. Specifically, we carefully design our prompts that include 1) task description, 2) travel characteristics, 3) individual attributes, and 4) guides of thinking with domain knowledge, and ask the LLMs to predict an individual’s travel behavior and explain the results. We select the travel mode choice task as a case study. Results show that, though no training samples are provided, LLM-based predictions have competitive accuracy and F1-score as canonical supervised learning methods such as multinomial logit, random forest, and neural networks. LLMs can also output reasons that support their prediction. However, though in most of the cases, the output explanations are reasonable, we still observe cases that violate logic or with hallucinations.
keywords:
Large language model; Travel behavior prediction
1 Introduction
Understanding human travel behavior has always been a crucial part of transportation planning. Travel behavior is a broad concept that consists of many attributes, such as the mode of transportation, the purpose of the travel, the choice of destination, and the time of departure. Consider travel mode choice as an example, which refers to how people make decisions on their means of travel. In the task of predicting mode choices, making accurate predictions is useful for both agencies and individuals. For agencies, it facilitates providing better transit services and forecasting travel conditions in the short term, while informing transportation infrastructure investments, better policies, and land use planning in the long run. In the traditional Four-Step Model for future demand and performance analysis, the mode share of a given population is estimated using mode split models (Ben-Akiva and Lerman, 1985). For individuals, on the other hand, mode choice prediction embedded in map applications helps them to make sound travel decisions. Conventional mode choice analysis relies on using numerical data to construct a mathematical model to represent human preferences (Mo et al., 2018, 2021). One example is the discrete choice model, where mode choice probabilities are calculated assuming people make choices to maximize their utilities (Ben-Akiva and Lerman, 1985).
Since last year, large language models (LLMs) have generated a tremendous amount of excitement in research due to their advanced language interaction and interpretation capabilities exhibited by-products such as ChatGPT. Having been trained on extremely large corpus and with parameters to the size of more human hundreds of billions, LLMs are able to interpret and execute diverse types of intricate tasks with a certain level of intelligence (Zhao et al., 2023). In transportation, studies have shown the abilities of GPT in making mobility predictions which outperforms state-of-the-art time-series machine learning models (Wang et al., 2023b). However, the question of whether LLMs can tackle the problem of travel behavior predictions is unclear.
In response to the above inquiry, this research aims to analyze the ability of LLMs to make travel behavior predictions. Specifically, this research focuses on travel mode choice prediction as a case study. Through prompt engineering, we are able to establish a zero-shot prompting framework that describes the task of the travel mode choice prediction, travel characteristics, and passenger attributes, as well as guiding the LLM with transportation domain knowledge. We demonstrate that although LLM-based predictions do not require any specific data inputs, they can achieve competitive performance than the classical models such as multinomial logit, random forest, and neural networks.
This paper makes the following contributions:
  • 1.
    This paper presents a framework to make travel behavior predictions with LLMs. To the best of the knowledge of the authors, this is the first effort to utilize the semantic power of LLMs in travel behavior prediction.
  • 2.
    The research demonstrates the accuracy and robustness of the framework by testing and evaluating it with real-world survey data and comparing the performance with classical benchmark models.
  • 3.
    The research proposes a paradigm shift in travel forecasting from using numerical data to establish mathematical models, to leveraging the reasoning ability of LLMs to predict travel behavior without training samples.
The remainder of this paper is organized as follows. The literature review is presented in Section 2. In Section 3, we describe the method for using LLM to predict travel behavior. Case studies and experiment results are shown in Section 4. Section 5 concludes the paper and discusses the future research directions.
2 Related work
2.1 Recent developments in large language models
Large Language Models (LLMs) typically refer to the language models based on the Transformer architecture with more than hundreds of billions of parameters, which have been shown to have abilities to excel in many complex tasks (Zhao et al., 2023). Wei et al. (2022) has demonstrated the emerging abilities of LLMs that only appear in larger models and cannot be predicted based on smaller models’ performance. Newer models such as GPT-4 illustrated strong capabilities in language processing and manipulation, quantitative reasoning, as well as planning and learning (Bubeck et al., 2023). Moreover, Gurnee and Tegmark (2023) has shown LLMs do have structured knowledge on both space and time using Llama-2 family models; they illustrated the model activations demonstrate linear representations of landmarks and historical events. In addition, products like ChatGPT which implement LLMs into chatbots and search engines that are easily accessible by the general public generated tremendous attention and excitement in this field (Zhao et al., 2023).
Apart from their powerful ability in complex tasks, LLM has brought a new research paradigm that leverages the interaction with prompt interfaces or APIs to develop applications as opposed to trained models from scratch (Zhao et al., 2023). The performances of these LLMs to user inputs often depend on the level of prompt engineering. Common methods to improve the performance of the tasks in LLMs include in-context learning (Brown et al., 2020; Dong et al., 2023), chain-of-thought (Wei et al., 2023) and plan-and-solve (Wang et al., 2023a; Zhou et al., 2023) to elicit the step-by-step reasoning power of LLMs (Zhao et al., 2023).
On the other hand, LLMs are also prone to several drawbacks that need to be considered when utilizing them. One limitation is Hallucination, where texts generated by LLMs could conflict with themselves or can not be verified (Bang et al., 2023). Another challenge of LLMs is their limit on recent knowledge since models trained on a big corpus of data with a large number of parameters typically cannot be updated in real-time (Zhao et al., 2023).
2.2 Large language model as a predictor
With the exceptional performance of LLMs, studies in many fields have utilized them to liberate traditional practices. ChatGPT has been shown to achieve high accuracies for linguistic-related classification tasks, such as genre recognition, personality prediction, political opinions, and hateful speech detection (Liu et al., 2023; Kuzman et al., 2023; Amin et al., 2023; Zhang et al., 2023; Huang et al., 2023). In healthcare, ChatGPT is used to facilitate tasks such as image annotations and finding summaries (Liu et al., 2023). For example, Caruccio et al. (2024) compared the performance of ChatGPT, other pre-trained language models, and traditional machine learning models on predicting low- and medium-risk diseases based on data on symptoms and diseases. The field of law also attracts many studies in LLM predictions due to its linguistic dependence. For instance, Trautmann et al. (2022) demonstrated the natural language processing ability of the GPT models by asking it to make legal judgment predictions from long legal documents.
In the field of transportation, several papers have shown the capability of LLMs to make predictions. Before the boom of LLMs, there have been several studies using pre-trained language models to predict mobility. Xue et al. (2022); Xue and Salim (2022) have been looking at converting mobility data into language descriptions, and using them as input data to fine-tune language models. With the rise of the GPT family, Wang et al. (2023b) demonstrated the powerful ability of GPTs to understand semantic sequential mobility data. In this study, each person’s historical locations, which are typically presented as structured numerical sequential data, are converted into language prompts and fed into the GPT to let it infer the next location. The predictions by LLMs outperform the traditional time-series models such as LSTM and the Multi-Head Self-Attentional (MHSA) neural network. Existing research offers promising potential for using LLMs as predictors. However, the existing literature only focuses on leveraging LLMs as a predictor in sequential mobility prediction and there has been no research exploring the ability of LLMs in travel behavior prediction tasks.
3 Methodology
3.1 Preliminaries
3.1.1 Large language model
Large language models (LLMs) are neural network-based architectures designed to process and generate human language text. Recent LLMs are usually built upon the transformer architecture. This architecture employs a deep neural network with self-attention mechanisms, allowing for the modeling of long-range dependencies in sequences of text. The size and scale of LLMs have grown exponentially, with models containing hundreds of millions or even billions of parameters. The increase in model size directly correlates with improved performance on various NLP tasks.
Pre-training is a crucial phase in the development of LLMs, where the model is exposed to massive amounts of text data to learn the statistical properties of language, including grammar, semantics, and world knowledge. The primary objective of pre-training is to train the model to predict the next word in a sentence or sequence of words. This task, known as language modeling, allows the model to capture the statistical regularities and contextual dependencies within the training data. Note that different LLMs (e.g., BERT and GPT) may have different pre-training tasks. In this study, we introduce the pre-training based on GPT’s framework for unsupervised multitask learning (Radford et al., 2019). The objective of pre-training can be expressed as:
$\boldsymbol{\theta^{*}}=\arg\max_{\boldsymbol{\theta}}\sum_{u\in\mathcal{U}}\sum_{i=1}^{n}\log\mathbb{P}\left(w_{i}^{(u)}\mid w_{1}^{(u)},w_{2}^{(u)},\ldots,w_{i-1}^{(u)};\boldsymbol{\theta}\right)$ (1)
where $\mathcal{U}$ is the set of all training corpus. $w_{i}^{(u)}$ is the $i$-th token of the $u$-th sequence. Given the trained parameter $\boldsymbol{\theta^{*}}$, we can use the model to generate answers for various tasks:
$s^{*}=\operatorname*{arg\,max}_{s}\mathbb{P}(s\mid\text{(Input, Task)};\boldsymbol{\theta^{*}})$ (2)
where $s^{*}$ is the output sequence with the largest probability (or relatively large probability depending on the searching algorithm and degree of randomness). $s^{*}$ is generated word by word until the “[End]” token is found. For example, asking the LLM to solve an addition problem can be expressed as
[…]
In the real-world implementation, the “[End]” token will be automatically added to the input sequence and will not be displayed at the end of the output sequence.
3.2 Conceptual framework
Given LLMs provide a generalized multitask solver, it is possible to use them as a predictor for people’s travel behavior prediction. The “(Input, Task)” in Eq. 2 is referred to as prompt in LLMs. The performance of LLMs for many tasks highly depends on the design of the prompt. We carefully design our prompts that include 1) task descriptions, 2) travel characteristics, 3) individual attributes, and 4) guides of thinking with domain knowledge. Details will be shown in Section 3.3.
The framework of the LLM-based travel behavior prediction is shown in Figure 1. Here we use the travel mode choice task as an illustration example. The input information will be organized and embedded into the prompt (i.e., prompt design). The prompt is fed into an LLM and outputs the prediction and associated reasons. It is worth noting that no supervised training step is needed.
Figure 1: Conceptual framework
3.3 Prompt design
Building upon existing prompting strategies, we carefully develop context-inclusive prompts that incorporate relevant contextual information to enhance travel behavior prediction. Detailed components are described as follows.
3.3.1 Task description
The task in this study is travel behavior prediction, which may include travel mode choice, travel purposes, travel departure time, duration, destination, etc. In the task description, we need to specify the available options and potential information we will provide. Example prompts are shown in Figure 2
Figure 2: Example prompts for task descriptions
3.3.2 Travel characteristics
People’s travel behavior is largely decided by travel characteristics, such as origin, destination, travel time, travel cost, etc. The prompts should include this information in a concise and organized way. In this study, we use a dictionary format the organize the travel characteristics information. Examples are shown in Figure 3.
Figure 3: Example prompts for including travel characteristics
3.3.3 Individual attributes
Individual attributes may also affect people’s travel behavior. This information is included in a descriptive way (instead of using dummy binary variables as typical mathematical models). Examples are shown in Figure 4.
Figure 4: Example prompts for individual attributes
3.3.4 Guide of thinking with domain knowledge
The ability of LLMs to perform complex reasoning can be improved by designing prompts with strategies like Chain-of-Thought (Wei et al., 2022) and Plan-to-Solve (Wang et al., 2023a). The essential idea is to guide LLMs on how to use the given information. This is the most important part of enabling LLMs to predict travel behavior without given any training data. The prompts should include domain knowledge with human common sense. The example of travel mode choice is shown in Figure 5.
Figure 5: Example prompts for the guide of thinking with domain knowledge
The first three aspects are domain knowledge. This can be customized based on input features and the specific task. The last paragraph is used to guide LLMs for numerical comparison. It is known that arithmetic and symbolic reasoning are challenging tasks for LLMs without well-designed prompts (Rae et al., 2021). Even for a simple task like providing LLMs with three numbers A, B, and C, and asking them to sort the numbers, LLMs can make many mistakes (this problem may be mitigated with more advanced LLMs). Therefore, we add the last paragraph, telling LLMs which mode has the lowest travel time or travel cost, and how large the relative differences are. Experiments show that this tends to be an effective way to enable LLMs for arithmetic reasoning in this context.
3.3.5 Interpretation and output
The last part of the prompt is to specify the outputs of the LLMs. The outputs include two parts: 1) the predictions and 2) the reasons that support the prediction. There are two motivations to ask for the reasons. First, it gives the model interpretability and helps us to diagnose the prompt design. Moreover, asking for explanations also improves LLMs’ reasoning ability. This is similar to the Plan-to-Solve prompting (Wang et al., 2023a), where the LLM’s reasoning performance is improved when the specific calculation steps are asked to output. The example prompt for travel behavior prediction is shown in Figure 6. Note that we also specify the output format so that we can easily extract the results.
Figure 6: Example prompts for interpretation and output
3.3.6 Summary
An example of the final prompt is shown in Figure 7. The sequence for different components is reorganized. It is worth noting that we do not include any training data information in the prompt (i.e., zero-shot), which allows the model to be used in any new context for travel behavior prediction.
Figure 7: Example complete prompts for travel behavior prediction (Colors before each sentence represent different components. They will not be input in the LLMs)
4 Case study
4.1 Data
The model is evaluated on the Swissmetro stated preference survey data set (Bierlaire et al., 2001). The survey aims to analyze the impact of travel modal innovation in transportation, represented by the Swissmetro, a revolutionary maglev underground system, against the usual transport modes represented by car and train. The data contains 1,004 individuals with 9,036 responses. Users are asked to select from three travel modes (train, car, and Swissmetro) given the corresponding travel attributes.
The LLM-based prediction model is compared against three benchmark models: 1) multinomial logit (MNL), 2) random forest (RF), and 3) neural networks (NNs). MNL is the canonical travel behavior prediction tool, while random forest and NNs are proven to perform well empirically (Wang et al., 2021). Since the sample sizes are small, we do not include more advanced deep-learning models as benchmarks.
Considering the cost of the OpenAI API, We further down-sample from the whole data set. We use a balanced sampling to randomly select 1,000 samples as the training data (only for benchmark models) and 200 samples as the testing data set. Balanced sampling means the proportions of each travel mode in the final data set are almost the same.
The features and data statistics (combined training and testing) are shown in Table 1.
Table 1: Training and testing data statistics
Features Mean Std Min Max
Train travel time (min) 160.6 74.0 34 794
Train cost (Swiss franc) 89.1 65.8 10 480
Car travel time (min) 139.1 98.0 32 1560
Car cost (Swiss franc) 87.9 49.4 10 520
Swissmetro travel time (min) 84.2 49.7 8 750
Swissmetro cost (Swiss franc) 110.2 83.5 14 768
Regular train user (Yes = 1) 0.34 0.47 0 1
Own train annual pass (Yes = 1) 0.13 0.34 0 1
4.2 Parameter settings
The specific LLM employed in the experiment is GPT-3.5 (version gpt-3.5-turbo-1106), which is one of the most advanced and widely used LLMs with open APIs. The prompts used are the same as Figure 7. We set the temperature to 0 to avoid randomness in the output. The benchmark models are implemented using the python sklearn package with the default hyper-parameters.
4.3 Results
4.3.1 Prediction performance
All models are evaluated based on prediction accuracy and weighted F1-score. Results are shown in Table 2. The LLM-based prediction has shown competitive prediction performance as many commonly-used supervised learning models, though we do not provide any training samples.
Table 2: Prediction performance comparisons
Models Accuracy F1-score
LLM 0.655 0.648
MNL 0.645 0.639
RF 0.650 0.646
NN 0.685 0.676
4.3.2 Interpretability
A unique advantage of LLM-based prediction is the interpretability of the results. Table 3 shows several example outputs from the LLM. In Case 1, LLM can correctly predict the travel mode choice (while MNL cannot) and provide reasonable explanations. In Case 2, though LLM makes the wrong prediction, the explanations are reasonable. It talks about the reason for choosing a Car due to its lowest travel cost, and its travel time is not too different from other modes. In Case 3, we observe LLM’s reasoning errors, stating that the absence of a train annual pass makes the person choose the train over Swissmetro. In Case 4, we see a common hallucination problem of LLMs (Bang et al., 2023). It claims that the person prefers time-saving over cost-saving without any related inputs.
Table 3: Example cases of using LLMs for travel mode choice prediction
Case 1: Correct prediction and reasonable explanations
Input:
{Travel time: {Train: 106, Car: 90, Swissmetro: 34}, Travel cost: {Train: 72, Car: 70, Swissmetro: 78}}. The person is not a regular Train user. He/She does not own the Train annual pass.
Prediction:
Swissmetro (MNL prediction is Car)
Actual:
Swissmetro
Reason:
Based on the given information, the person is not a regular Train user and does not own the Train annual pass. Swissmetro offers the lowest travel time, saving 62% compared to Car and 68% compared to Train. Although Car has the lowest travel cost, the significant time saving of Swissmetro makes it the most likely choice for the person.
Case 2: Wrong prediction but reasonable explanations
Input:
{Travel time: {Train: 114, Car: 96, Swissmetro: 55}, Travel cost: {Train: 82, Car: 80, Swissmetro: 105}}. The person is not a regular Train user. He/She does not own the Train annual pass.
Prediction:
Car (MNL prediction is Car)
Actual:
Train
Reason:
Based on the given information, the person is not a regular Train user and does not own the Train annual pass. Considering the trade-off between travel time and cost, the person is likely to choose the Car as it has the lowest travel cost and the difference in travel time compared to Swissmetro is not significant enough to outweigh the cost savings.
Case 3: Wrong prediction and unreasonable explanations
Input:
{Travel time: {Train: 124, Car: 96, Swissmetro: 59}, Travel cost: {Train: 83, Car: 80, Swissmetro: 133}}. The person is a regular Train user. He/She does not own the Train annual pass.
Prediction:
Train (MNL prediction is Train)
Actual:
Swissmetro
Reason:
The person is a regular Train user, and regular Train users prefer to use Train. Although Swissmetro has the lowest travel time, the person’s preference for Train and the absence of a Train annual pass make it more likely for the person to choose Train over Swissmetro. Additionally, Car has the lowest travel cost, but the person’s regular use of Train and preference for it make it more likely for the person to choose Train over Car despite the cost difference.
Case 4: Wrong prediction and hallucinated explanations
Input:
{Travel time: {Train: 95, Car: 130, Swissmetro: 92}, Travel cost: {Train: 29, Car: 44, Swissmetro: 32}}. The person is not a regular Train user. He/She does not own the Train annual pass.
Prediction:
Swissmetro (MNL prediction is Train)
Actual:
Train
Reason:
Based on the given information, the person is not a regular Train user and does not own the Train annual pass. Swissmetro has the lowest travel time, saving 29% compared to Car and 3% compared to Train. Although Train has the lowest travel cost, the person’s attributes suggest a preference for time saving, making Swissmetro the most likely choice.
5 Conclusion and future work
In this study, we show the potential of using LLM for the complex travel behavior prediction task without using training samples. The LLM-based prediction has competitive performance as many commonly used supervised learning models such as MNL, RF, and NNs. Besides, LLM can also output explanations for the inference, though there are still reasoning errors and hallucination problems.
This paper presents a preliminary result, there is a lot of ongoing and future work. 1) First, more experiments are needed, including using larger data sets, conducting replicated experiments to mitigate random errors, and replicating the study in other travel behavior tasks such as travel purpose prediction. 2) The proposed prompt is composed of many components. Ablation studies can be conducted to test the effectiveness of different components in the prompt design. 3) The current prompt is a zero-shot prompt. Future studies can include some representative samples in the training data set as a few-shot prompt for in-context learning. The few-shot prompt needs to guide the LLMs to learn to quantify the value of time (or other important concepts for prediction) using the limited samples. 4) More experiments can be conducted on testing the performance of different LLMs as well.